if x varies inversley as v, and x=45 when v=2, find x when v=10.
Answer by uma(370)

x varies inversely as v
==> x = K/v for some constant K
given x = 45 when v = 2.
Plugging in these values in the above ,
45 = K/2
==> K = 90
So x = 90/v
When v = 10,
x = 90/v
= 90/10
= 9
So when v = 10, x = 9